OpenStudy (anonymous):

who can give me the answer to this question If g(x) = 3x + 1, evaluate g(-14).

OpenStudy (anonymous):

just put -14 in place of x in g(x) = 3x + 1

OpenStudy (anonymous):

g(x)= 3x+1 g(-14) = 3(-14)+1 =-42+1 =-41